In y=4x, what happens to y when x is doubled?increased by 2?decreased by 3?halved?
Answer by KnightOwlTutor(293)

When x is doubled we write it as 2x and replace the x value with 2x
y=4(2x) y=8x The y value is also doubled 8x/4x=2
When x is increased by 2
y=4(x+2) y=4x+8 The y value is increase by 8 units
When x is decreased by 3
y=4(x-3) y=4x-12
The y value decreases by 12 units
When x is halved x/2
y=4(x/2)
y=2x The y values are also halved.